Shift report 20011230 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Dec 30 2001 - 00:04:11 EST

  • Next message: Apache: "Shift report 20011230 00:00-08:00"

    ------------------------------------------------------------
    Date: Sun Dec 30 00:04:11 2001, Shift: 16:00-24:00
    Supervisor: R. Debbe, with: 
    Log book#11 pages 117 - 119
    
    Shift Summary:
    ------------------------------------------------------------
    1. When I arrived an access was scheduled for Phenix in the next half hour. 
    it took 6 hours to get beam back, and after a mere half an hour of high intensity 
    unpolarized store, the beam was dumped by a beam loss monitor.
    
    2. I only collected 25K trigger 6 events in run 6201
    
    3. During shift check I found that according to brahvo, 
    T2 drift voltage  was -192V eventhough the current was 
    -59.4 micro amps. The data from the day shift looks good.
    I decided to go in the IR and switch the channels in the L1458A
    prower supply. Now the voltage readout is correct. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Dec 30 2001 - 00:04:50 EST